OpenStudy (taepoohwae):

Can some please show me how to write this in standard form? (4 x 10^2)(1 x 10^3)(5 x 10^-4)

jimthompson5910 (jim_thompson5910):

the bases are: 4, 1, 5 multiply them out to get 4*1*5 = 20 the exponents are: 2, 3, -4 add them to get: 2+3+(-4) = 1

jimthompson5910 (jim_thompson5910):

so (4 x 10^2)(1 x 10^3)(5 x 10^-4) = 20 x 10^1 you can optionally convert this to scientific notation if you want, but ultimately you'll need to convert to standard form

OpenStudy (taepoohwae):

So how is tis standard form

jimthompson5910 (jim_thompson5910):

it's not, you need to convert 20 x 10^1 to standard form

OpenStudy (taepoohwae):

I think I understand what you're saying

jimthompson5910 (jim_thompson5910):

ok and what do you get for your final answer?

OpenStudy (taepoohwae):

200

jimthompson5910 (jim_thompson5910):

good, you got it

OpenStudy (taepoohwae):

Thanks a bunch!

jimthompson5910 (jim_thompson5910):

you're welcome